### 24 Mar 2026 v2

# The search for Population III: Confirmation of a HeII emitter with no metal lines at z=10.6

Abstract
We report the confirmation of a HeII1640 emitter located at 3 pkpc from the galaxy GN-z11, at z=10.6. The detection, based on JWST NIRSpec-IFU high-resolution spectroscopy, confirms a previous claim based on medium-resolution spectroscopy. The HeII1640 identification is further supported by the independent detection of H obtained by Übler et al. (2026) at the same location. The HeII emission is spectrally resolved in two components separated by 120 km/s. The Equivalent Width of the HeII emission is extremely high (20Å). No metal lines are detected. Population III stars appear to be the most plausible explanation for the observed HeII emission. We also discuss the possible contribution from a Direct Collapse Black Hole, or a Primordial Black Hole — these scenarios are less plausible, but cannot be ruled out completely.
keywords:
galaxies: high-redshift – galaxies: evolution – galaxies: abundances – stars: Population III – infrared: galaxies1 Introduction
The detection and characterization of the first population of stars, formed out of pristine clouds of gas, the so-called Population III (PopIII) stars, is one of the most important goals of modern astrophysics. These stars are theorised to form in galactic embryos within the first few hundred million years after the Big Bang (e.g., Abel et al., 2002; Yoshida et al., 2003; Klessen and Glover, 2023). There have been detections of galaxies with extremely low metallicities, in the range (e.g., Vanzella et al., 2023; Fujimoto et al., 2025; Hsiao et al., 2025; Maiolino et al., 2025; Morishita et al., 2025; Nakajima et al., 2025). These could be very low metallicity PopII galaxies, or possibly self-polluted or hybrid PopIII galaxies (Rusta et al., 2025). However, these metallicities are still well above the expected threshold for PopIII, which are generally identified below a critical metallicity , which prevents the formation of low-mass stars because of inefficient gas cooling (e.g., Bromm et al., 2002; Schneider et al., 2002; Bromm and Yoshida, 2011).
An alternative approach is to identify other tracers of PopIII. Most models expect PopIII stars to be very massive and very hot (e.g., Trussler et al., 2023; Schaerer et al., 2025; Storck et al., 2025; Wasserman et al., 2026a), and their hard spectral energy distribution (SED) is expected to doubly ionize helium, resulting in significant emission of HeII recombination lines, such as and (e.g., Schaerer, 2003; Zackrisson et al., 2011; Nakajima and Maiolino, 2022; Lecroq et al., 2025; Schaerer et al., 2025; Storck et al., 2025; Wasserman et al., 2026a).
While most of the PopIII stars are expected to form in small haloes, various models have predicted that they may also form around massive haloes at high redshift as a consequence of gas accretion, inefficient mixing, and massive star formation resulting from gas compression and Lyman-Werner heating from the central source (e.g., Liu and Bromm, 2020; Venditti et al., 2023, 2024, 2025). Interestingly, Venditti et al. (2023) show that it is potentially more likely to find PopIII stars in distant massive haloes than in their low-mass counterparts. While massive haloes are considerably rarer than their low mass counterparts in the early Universe (Wechsler and Tinker, 2018), they typically host massive and luminous galaxies, which makes them easier to identify in observations.
Within this context, Maiolino et al. (2024b) explored the immediate environment around the most UV-luminous galaxy at , GN-z11 (RA 12:36:25.44, DEC +62:14:31.3 Oesch et al., 2016; Bunker et al., 2023; Álvarez-Márquez et al., 2025). This galaxy has a stellar mass of , is located in an overdense region (Tacchella et al., 2023; Scholtz et al., 2024b), and likely hosts an AGN (e.g., Maiolino et al., 2024a; Ji et al., 2025; Scholtz et al., 2025b; Crespo Gómez et al., 2026; Fabian et al., 2026). Maiolino et al. (2024b) used NIRSpec-IFU medium resolution () spectroscopy to confirm a previous tentative detection of HeII1640 identified with the Prism (), along the MSA shutter targeting GN-z11. Specifically, they identified an emission line in a region located at about 3 kpc NE of GN-z11, at a wavelength consistent with HeII at a redshift very close to GN-z11 (). Given the inferred high equivalent width (EW) of the line (20Å) and the absence of metal lines, Maiolino et al. (2024b) suggested this to be a star-forming clump with possible signatures of PopIII stars.
In this paper, we report deep high-resolution (G235H; ) NIRSpec IFU observations of the same region targeted in the previous study. These new observations confirm the presence of HeII emission in the halo of GN-z11. Further confirmation is obtained through the detection of H at the same location and same redshift, which is reported in a companion paper (Übler et al., 2026).
A second companion paper, Rusta et al. (2026), presents a theoretical interpretation of our findings on this source in the Pop III scenario, also discussing the implications for the still-unknown mass distribution of Pop III stars.
Throughout this paper, and in the companion papers, this source is named Hebe111HElium Balmer Emitter. In ancient Greek mythology Hebe (H) is the goddess of youth, daughter of Zeus and Hera..
A flat CDM cosmology is adopted throughout based on the latest results of the Planck collaboration (Planck Collaboration et al., 2020), with , , .
2 Observations and data processing
Observations were obtained through the JWST programme ID (PI: Maiolino) using the NIRSpec IFU (Jakobsen et al., 2022; Böker et al., 2022, 2023) with the high resolution grating G235H (R2700), covering a field of view of . The observations were performed between May 15 and 17, 2025, split into two visits. Detectors were read out with the NRSIRS2 pattern, with 25 groups and 2 integrations. A total of 39 dithers were executed, with a medium cycling pattern. Unfortunately, during dither position 18 the telescope guiding camera locked onto a hot pixel; hence, this frame had to be discarded in further processing. After removing this frame, the total exposure time was 38.8 hours. The data reduction uses the JWST Science calibration pipeline v1.15.0 with CRDS jwst_1293.pmap. To improve the quality of the datacube, we performed a number of further data processing steps, in addition to the default pipeline. The customised procedures used for residual cosmic rays snowballs flagging, subtraction of pink noise (1/f noise), and removal for failed open MSA shutters signal are described in detail in Perna et al. (2023). Stage 2 was executed using the bad-pixel self-calibration implemented in the standard pipeline, in which all exposures on a given detector are used to identify and flag bad pixels that may have been missed by the reference bad-pixel mask. For the final stage of the pipeline, outlier rejection was performed following D’Eugenio et al. (2024b), with a rejection threshold at the percentile. Before combination, the cube was resampled to a scale of per spaxel.
The error extension of the cube is known to underestimate the actual error. Therefore, following Übler et al. (2023), we have rescaled the error extension to match the measured noise in regions of the cube free of sources.
The cube was registered with the NIRCam images by producing pseudo broad-band maps, by collapsing the cube across the large wavelength ranges matching three NIRCam filters. Unfortunately, only two sources in the field are available, GN-z11 and a relatively large foreground galaxy at . This, in addition to the coarse pixel size of NIRSpec, results in an accuracy of the resulting registration of about 0.05′′.
3 Hebe: confirmation of a HeII emitter in the halo of GN-z11
The upper panel of Fig.1 shows the spectrum extracted from a aperture centred at East and North of GN-z11 (RA 12:36:25.520, DEC +62:14:31.934). The spectrum shows an emission feature, detected at 6, which is identified as HeII1640 at (centroid of the line), blue-shifted by about 450 km/s relative to GN-z11. Note that, for this paper, the redshift of GN-z11 is consistently remeasured based on its HeII emission in our cube, and found to be ; this is slightly lower than the redshift given by Bunker et al. (2023) using medium resolution spectroscopy - - but it is also slightly higher than the redshift that would be derived from their HeII medium resolution detection, which would give ; these slight inconsistencies are likely due to the lower spectral resolution of the previous observations.
This independently confirms the previous HeII detection at the same location obtained with medium resolution spectroscopy by Maiolino et al. (2024b).
As already mentioned, we name Hebe this HeII emitting source, which is located at 3 kpc from GN-z11.
Maiolino et al. (2024b) already pointed out that no other line identification was plausible, because of the absence of other nebular lines, and that the HeII1640 line at a redshift very close to GN-z11 was the most likely explanation. Any possible ambiguity with potential lower redshift interlopers is now completely removed, as the HeII identification is unambiguously and independently confirmed by the detection of H, at the same location and redshift of Hebe, as reported in a companion paper by Übler et al. (2026). The H detection is shown for convenience in the bottom panel of Fig.1, while its thorough analysis is discussed in Übler et al. (2026).
Fig.2 shows the HeII map by collapsing the continuum-subtracted spectral channels between 1.8995 m and 1.9007 m (the three central channels of the line, i.e. a velocity interval of 187 km/s), which highlights the location of Hebe. The white circle indicates the location of GN-z11. Fig.3 shows the HeII contours again, but overlaid on the NIRCam RGB image, with the Field of View adjusted to show the location of Hebe relative to GN-z11 and a foreground galaxy at z=2.04, which causes some gravitational lensing (see Appendix C).
As discussed in more detail in Section 6 below, the location and flux of the HeII emission are broadly consistent with the previous detection based on the NIRSpec medium resolution grating (Maiolino et al., 2024b), although the comparison with the latter suggests the presence of some diffuse emission, both spatially and in velocity. Fig.2 shows the compactness of the line emission, which is unresolved ( 400 pc in size) at the NIRSpec resolution. We note some diffuse emission extending towards the SW. Even though its surface brightness is not statistically significant (mostly within of the background fluctuations), a similar diffuse emission was seen in the medium resolution map. This will be discussed again in Section 6. There is also another marginally detected emitter 0.8′′ North of GN-z11; this was also marginally detected in medium resolution map, and will be discussed in Section 6.
The HeII emission has a FWHM of about 110 km/s (deconvolved by the spectral resolution)222Following, e.g., D’Eugenio et al., 2025 or Shajib et al., 2025, we take the spectral resolution as the nominal value (Jakobsen et al., 2022), but at 1.9 m this is divided by a factor of 0.7 (determined empirically in those references), which takes into account that, for compact sources, the size of the source can be comparable to the size of the IFU slices., but this apparent width is actually resulting from two marginally resolved components. This is in contrast with the H emission, which is spectrally unresolved (see Übler et al. 2026; note that the two spectra have similar resolution – at 1.9m and at 5m). We have simultaneously fitted HeII and H with two Gaussian components (C1 and C2), as shown by the dashed lines in Fig.1. The two components are separated by km/s, and each of them is consistent with being spectrally unresolved (. The simultaneous fit indicates that H is associated only with C2, while the component C1 of H is undetected.
We have also attempted a simultaneous fit with a single Gaussian component, which results in a significantly worse fit, as shown in Appendix A.
It is difficult to explore whether there is any small spatial offset between the two components C1 and C2, as the maps of individual spectral channels become too noisy for a proper assessment.
The inferred fluxes and luminosities of the HeII line and its two individual components are reported in Table 1. These include an aperture correction factor of 3.03, estimated from the cubes of two stars (ID 1219 –PI: Luetzgendorf – and ID 1222 –PI: Willott–) processed with the same methodology, at the wavelength of 1.9m. The total flux has been estimated by simply summing up the flux in the central 5 spectral pixels of the line. The flux of the individual components is determined by the simultaneous double Gaussian fit.
We also correct the luminosities for a magnification of 1.42 due to the foreground galaxy at z=2.028. Details of the lensing model and magnification estimation are given in Appendix C.
As discussed in Section 6, there is evidence for a diffuse component of HeII; in the second part of Table 1 we also provide an estimate of Hebe’s HeII flux after subtracting this more diffuse component333This is done by taking the diffuse flux estimated in Sect.6 and deriving the contribution to the extraction aperture..
| Aperture corrected fluxes | |
| Total | |
| Å | |
| <3.1 (<2.7)b nJy | |
| Component 1 | |
| Å | |
| Component 2 | |
| Å | |
| De-lensed luminosities and cleaned of diffuse component | |
4 Equivalent width constraints
Measuring or constraining the equivalent width of the HeII emission is incredibly important, as it can provide key constraints on the nature of the source. This requires a detection of the continuum emission. For extremely low metallicity galaxies, the main difficulty is the continuum detection rather than the nebular line emission. A clear example is the extremely metal poor galaxy LAP-1 at (Vanzella et al., 2023), where multiple nebular emission lines are clearly detected, while the continuum remains undetected despite deep exposures and lensing magnification (Nakajima et al., 2025). The HeII emitter in GN-z11 has a similar issue, with the additional complication that it is in the proximity of the relatively extended foreground galaxy at z=2.028. As discussed in Maiolino et al. (2024b) (see also Tacchella et al., 2023), modelling and removing the emission from the foreground galaxy leaves some residual flux around the location of Hebe. Yet, there is no clear indication of a continuum point source associated with it, suggesting that the flux is some diffuse residual after the foreground galaxy removal. Thanks to the more precise localization of Hebe, we can now confirm the absence of a continuum detection at the 5 upper limit level of 3.1 nJy in F200W. We derive this value from the same aperture adopted for the spectrum extraction, accounting for a correction for a point-like source. Direct photometry (also corrected for the limited aperture size) provides a flux measurement of nJy. The 5 upper limit was corrected for the contribution of the HeII line to the broad-band measurement to get a spectral continuum upper limit. The rest-frame EW(HeII) corresponding to this continuum is 47 Å. We provide a conservative lower limit on the EW of the C1 and C2 components by applying the same continuum upper limit to each of them. This translates into lower limits of 28 Å and 22 Å, for the C1 ad C2 components, respectively. All these measurements are provided in Table 1.
As we discuss extensively in Section 7, the inferred rest-frame EW(HeII) are higher than 20Å, well above the EW expected for a metal-enriched (PopII) stellar population. If the EW(HeII) were Å then we would have expected the detection of a source with a flux higher than 6 nJy in F200W.
5 Absence of metal emission lines
The G235H/F170LP setup covers a wavelength range where various metal emission lines typically detected in high-redshift galaxies (e.g. D’Eugenio et al., 2024a; Scholtz et al., 2025a) are expected to be seen at the redshift of Hebe. Specifically, the NIV]1483,1486, CIV1548,1551, OIII]1661,1666, and CIII]1906,1908 doublets, as well as the NIII]1749 multiplet, are all within the spectroscopic range of this setup. Considering that in local and high-z galaxies at least some these lines (if not all of them) are stronger than HeII, at least some of these metal lines should be visible also in Hebe. On the contrary, none of these transitions is detected in Hebe. We report conservative upper limits of the flux of these metal lines relative to HeII. The spectral regions around the location of these lines are shown in Appendix A.
Moreover, in the companion paper by Übler et al. (2026) we show that no lines are detected in the G395H spectrum aside from the prominent H (and a tentative detection of H). In particular, we do not detect [NeIII]3869, whose strength is often comparable to H in high-z galaxies (Scholtz et al., 2025a).
The companion paper by Rusta et al. (2026) provides a theoretical interpretation of our observational findings, showing that these non-detections are very constraining of the nature of the ionizing sources.
6 Consistency with the medium resolution observation and diffuse HeII emission
In this section, we compare the HeII detection in our new high-resolution (G235H) data to the previous detection in the medium resolution spectrum (G235M) presented in Maiolino et al. (2024b).
The aperture-corrected HeII flux of Hebe (and not corrected for magnification) is . This is to be compared with the flux of reported by Maiolino et al. (2024b). The two fluxes, therefore, are marginally consistent, within . However, the comparison is more complex, as discussed in the following.
To begin with, we compare the HeII spectra of GN-z11 in the two cubes and notice that there is probably a slight wavelength miscalibration between the two, in the sense that the G235M spectrum requires a blueshift by about 0.6 spectral pixels to be aligned with the G235H spectrum. It is not clear whether this is due to a potential issue in the calibration, or to different versions of the calibration files. However, investigating these aspects is beyond the scope of this paper, and we simply apply the slight wavelength shift to the G235M spectrum.
Secondly, in Fig.4 we compare the R2700 HeII map obtained in this work, with the R1000 HeII map obtained in Maiolino et al. (2024b) (contours). Note that the G235M map shows significant HeII emission associated with GN-z11, not seen in the G235H map because the velocity range integrated in G235M is larger and includes also some HeII emission from GN-z11 (recall that Hebe is blueshifted by km/s relative to GN-z11). We have astrometrically registered the two images via the GN-z11 continuum. Unfortunately, this leaves some uncertainty in the relative positions. Specifically, in addition to the coarseness of the NIRSpec spatial sampling ( native spaxels), there are only two continuum sources that can be used for the alignment: GN-z11 and the foreground galaxy, which is not point-like but fairly extended. The white circle in the figure illustrates the location of GN-z11, and its size illustrates the uncertainty in aligning the two continua. The HeII emission in G235H appears slightly offset towards the NE relative to the emission seen in G235M, although within the maps relative alignment uncertainty. It is also true that the G235M emission is not symmetrically distributed and its peak corresponds to the maximum emission in G235H, within half a (resampled) pixel. It is possible that there is a residual alignment issue, but it is also possible that the two maps do not match exactly because they are probing slightly different velocity ranges, as discussed below.
The HeII flux reported in Maiolino et al. (2024b), was integrated on a larger velocity range (simply because of the lower resolution and larger spectral channels), and also in a larger aperture (), centred more towards the South relative to the smaller () aperture used in this paper. We therefore re-extract a spectrum on an aperture and velocity range matching the extraction of Maiolino et al. (2024b) – the spectrum is much noisier, but the total flux is , hence matching very well the flux observed in G235M. This indicates that there is some extended HeII flux, diffuse both spatially and in velocity, that is probably not clearly visible in the high-resolution data due to the higher noise affecting the signal distributed on larger velocities. Some of this diffuse emission is actually potentially detected in the HeII maps of Figs.2 and 4, and was potentially seen also in the medium resolution map of Maiolino et al. (2024b). The origin of such diffuse emission, probably with high velocity (of the order of a few hundreds km/s), is not clear. However, we note that Hamel-Bravo et al. (2025) detected broad (655 km/s) HeII emission (without corresponding broad Balmer emission) in the halo of the ultra-metal poor local galaxy SBS 0335-052E; the nature of such broad HeII emission is uncertain, but could result from near-pristine gas shocked to high temperature by the outflow driven by the central galaxy. A similar scenario could be happening in the vicinity of Hebe: the vigorous star-forming activity may have driven an outflow that is shock-heating the gas to high temperatures. However, the inferred velocities are quite modest ( km/s), so it is unlikely that shocks are relevant (Lecroq et al., 2024). Additionally, we warn that the presence of high-velocity gas is inferred only indirectly from the total flux in the large aperture, and when compared with the medium-resolution data, deeper data would be needed to directly confirm it.
We finally note that some diffuse emission was identified in the NIRCam images in the vicinity of GN-z11, within a radius of about North of GN-z11, the so-called ‘Haze’ (Tacchella et al., 2023). This is marginally visible in Fig.3. The origin of this emission is not clear, but Tacchella et al. (2023) suggests that it is a foreground, low-redshift interloper. Here we simply notice that such ‘Haze’ does not seem to be physically associated with the tentative diffuse emission seen in HeII.
7 The Population III scenario
Spectra with prominent HeII emission and absence of metal lines are expected in the case of Population III stars (e.g. Zackrisson et al., 2011; Trussler et al., 2023). In this section, we explore this scenario more quantitatively through a simple comparison with some models and expectations.
Fig.5 shows the EW(HeII1640) vs HeII/H diagnostic diagram. This diagram was already presented in Maiolino et al. (2024b) for Hebe. In that paper, however, H was measured along the shutter using the prism, hence likely only grazing at the more diffuse emission and not probing the core of the Hebe (see discussion in Section 6). With the new data presented in Übler et al. (2026), we have a direct measurement of H in the same region and with a similar spectral resolution as for the HeII emission. The various symbols show models from Nakajima and Maiolino (2022), with various ionization parameters and metallicities. Given the complete absence of metal lines, to avoid overcrowding the plot, we focus on low metallicity models, below 10% solar. Specifically, red stars indicate PopII models, with decreasing metallicity from dark to light, down to a metallicity of (i.e. ). Note that these models use BPASS binary stellar templates, which are already more energetic than standard stellar populations (see Nakajima and Maiolino, 2022, for details). Orange pentagons are AGN models, with a broad range of possible ionizing spectra, a broad range of ionizing parameters, and a very broad range of metallicities, from 10% solar to completely pristine. The squares show the PopIII models, with sizes indicating the adopted IMF (assuming a Salpeter slope): larger symbols are for stellar masses in the range , intermediate size symbols are for stellar masses in the range , and small symbols for stellar masses in the range . The blue squares are for gas densities of . The purple squares are for gas densities of and . The small blue points are PopIII models from Rusta et al. (2026): dark points are for Pure III pristine systems, while light points are for systems with some PopII contamination, but still dominated by PopIII (50%). The blue points are for gas densities of and ionization parameter , the purple points are for gas densities of and ionization parameter .
We note that densities of explored in the models may appear somewhat extreme; however, high redshift studies have revealed a significant population of galaxies with gas densities in this range (e.g. Martinez et al., 2025; Moreschini et al., 2026). Therefore, it is very plausible that PopIII systems can have such high densities. The reason why PopIII models with high densities are significantly higher in EW is that these are above the critical density of the 2-photons nebular continuum, which is therefore suppressed, boosting the EW(HeII). Interestingly, the same does not happen for AGN; indeed, in these cases, the 2-photon continuum is negligible, and the UV continuum is dominated by the accretion disc. Therefore, increasing the gas density to for the AGN models does not have any tangible effect.
The large golden symbols are the values measured for Hebe, both total (triangle) and for the individual components (circles), as labelled. The EW(HeII) and HeII/H ratios observed for Hebe, either total or for the individual components, are well above the expectations for PopII models. AGN models also struggle to reproduce the observed constraints, primarily because they generally predict lower EW(HeII); however, we will discuss in more detail the AGN scenario in section 10. Component C1 is the one most clearly aligned with the PopIII prediction, both for the low- and high-density models. Component C2 seems inconsistent with PopIII low-density models, but still consistent with PopIII high-density models. We shall mention that it is possible that other PopIII models, with steeper SED and higher yield of HeII photons, which for example are associated to rotating or binary PopIII stars (e.g. Yoon et al., 2012; Murphy et al., 2021; Sibony et al., 2022; Wasserman et al., 2026b; Lecroq et al., 2025), could reproduce C2 with low density gas. These models were not considered in Nakajima and Maiolino (2022) nor in Rusta et al. (2026). An alternative possibility is that in C2 the very first generation of PopIII may have provided a first quick enrichment of dust, as expected by many models (see review by Schneider and Maiolino, 2024). Taking the extinction curve by Sun et al. (2026) (one of the few extinction curves measured at high redshift), an extinction of only would be enough to move the HeII/H flux ratio to the PopIII sequence. Depending on the dust distribution geometry, in a compact configuration such an extinction can be achieved with just a few solar masses of dust (see e.g. Madau and Maiolino, 2026).
Overall, both for C1 and C2 excitation by PopIII stars appears to be the scenario most consistent with the current observational constraints on Hebe. In this scenario, the two components are two nearby clusters (within 400 pc of each other), probably at slightly different evolutionary stages (possibly with C2 being already contributed by some PopII) and/or slightly different local physical properties (different densities and ionization parameters).
A more extensive comparison with more detailed models is provided in the companion paper by Rusta et al. (2026), who reach similar and even more stringent conclusions. They also constrain the total stellar mass of the PopIII star clusters in the range between and a few times . We note that is already well in the range expected for PopIII systems (e.g. Storck et al., 2025), although possibly towards the upper end, which is not unexpected given the peculiar environment. One should also take into account that there are various uncertain factors that could result in further reducing the inferred stellar mass, specifically: i) the lensing factor of the foreground galaxy is uncertain, and could result in a mass two times lower; ii) as mentioned, other PopIII models (especially rotating and/or binary PopIII stars) can result higher luminosity for a given mass, which would alos result in a reduced mass; iii) as we have spectrally resolved Hebe in two components, it is likely that also these two components are made of even lower mass systems, whose individual spectral signatures are blended within our spectral resolution.
We finally discuss that, as already mentioned in Section 3, there is another secondary peak in the G235H HeII map, located at about 0.8′′ North of GN-z11. Based on the R2700 alone, this would certainly not be significant. However, it is interesting that it is marginally detected at the same location also in the G235M map, which may support that this is a second HeII emitter. If so, this would fit a scenario in which the halo of GN-z11 has recently been subject to pristine gas accretion from the N-NE, which is resulting in PopIII star formation at various locations as a consequence of gas compression. We note that there is possibly a third peak to the NW of GN-z11 seen both in the G235H and G235M maps; however, this is weaker in G235H and is also located in the region of G235M (West of GN-z11) where there are fewer frames and possible edge effects, hence less reliable.
8 Lyman-
Most models of PopIII stars expect these to power strong Ly emission (Raiter et al., 2010). However, at the IGM absorption of Ly is extremely strong, so that the vast majority of galaxies do not show any Ly emission, despite often displaying prominent nebular emission lines (e.g. Tang et al., 2024; Jones et al., 2024; Kageura et al., 2025; Boyett et al., 2024). There are a few rare cases of Ly emitters at high redshift (Ono et al., 2012; Finkelstein et al., 2013; Oesch et al., 2015; Zitrin et al., 2015; Roberts-Borsani et al., 2016; Larson et al., 2022; Tang et al., 2023; Cooper et al., 2024; Witstok et al., 2025b; Leonova et al., 2025), the most remarkable one being JADES-GS-z13-1-LA at (Witstok et al., 2025a). This galaxy has likely carved a small ionized ‘bubble’ within the nearby IGM that allows its Ly to escape. Despite its very high redshift, GN-z11 also shows the presence of some weak and extended Ly emission (Bunker et al., 2023; Scholtz et al., 2024a), although the presence of an ionised bubble is not strictly required. The Ly emission in GN-z11 is significantly redshifted (by ), and much weaker than expected from the Balmer lines (; Bunker et al., 2023), indicating that it is in fact heavily suppressed by IGM damping-wing absorption.
An IFU G140M observation of GN-z11 covering the location of Hebe is available (Scholtz et al., 2024a). This observation is much shorter than those with G235M or G235H. Ly is not detected at the location of Hebe, with a upper limit of . This includes the aperture correction and is integrated over a velocity range of 150 km/s. Also, H is 5 times weaker in Hebe compared to GN-z11, suggesting that the intrinsic Ly luminosity of this regions is lower by a comparable factor (PopIII and AGN are expected to have a similar HeII/Ly intrinsic ratio, Nakajima and Maiolino, 2022).
We have simulated IGM transmission curves at , the systemic redshift measured by Übler et al. (2026), using the Mason and Gronke (2020) model adopted in the lymana_absorption code (Witstok et al., 2025b, a)444Publicly available at https://github.com/joriswitstok/lymana_absorption.. We conservatively assume an intrinsic line FWHM of , almost twice that of H. We consider the IGM to be fully neutral at this redshift, since for GN-z11 itself this correctly reproduces the inferred Ly transmission () at the observed velocity offset of the line (Bunker et al., 2023). As shown in Figure 6, for velocity offsets below , appropriate for a faint, low-mass system (; e.g. Mason et al., 2018), the predicted overall Ly transmission is well below , bringing the expected observed flux in agreement with the upper limit. Specifically, the expected Ly flux would be , i.e. more than an order of magnitude below the IFU upper limit.
The observed upper limit on Ly would only be violated if the velocity offset were extremely high, , relative to the system rest frame (i.e. the Ly tail extending even further redwards than what is seen in GN-z11). Additionally, we note that if Hebe is truly a satellite located within of GN-z11 as suggested by their close on-sky separation (), the shift of Hebe relative to GN-z11 represents a peculiar velocity (rather than a cosmological redshifting effect, which would correspond to a line-of-sight separation of ). Compared to GN-z11, whose Ly emission is already highly attenuated, the implied (additional) relative motion towards the neutral IGM along the line of sight would result in an even further suppression of the Ly emission from Hebe (dashed black line in Figure 6).
Therefore, Ly is essentially impossible to be detectable from Hebe, and this holds even for future, ultra-deep observations (of the order of 100 hours). The negligible Ly flux expected to reach us is thus fully consistent with our upper limit.
9 The Wolf-Rayet Scenario
One alternative interpretation of the observed HeII emission is the presence of a population of Wolf-Rayet (WR) stars. Specifically, in this scenario the broader HeII line, if not interpreted as two separate components, could be tracing a population of very metal-poor WR stars. Although broader than H, the HeII line is much narrower (only 110 km/s) than typically observed in WR stars. Yet, the width of the HeII lines in the WR expanding envelopes is expected to decrease as a function of metallicity, as most of the radiation pressure generating the envelope winds happens via metal line-driven winds. It is indeed observed that WR stars in more metal-poor environments tend to have narrower HeII (Vink and de Koter, 2005; Crowther and Hadfield, 2006; Crowther et al., 2023; Castro et al., 2018). However, even these low metallicity cases present prominent nitrogen (or carbon) emission lines, as (by definition of WR) the expulsion of the outer layers exposes the inner He and N burning core (e.g. Hainich et al., 2014; Berg et al., 2024, 2025).
The WR phenomenon in the extremely metal-poor regime (), has been studied only theoretically. The photospheric wind velocities are expected to decrease steadily as a function of metallicity (Gräfener and Hamann, 2008; Sabhahit et al., 2023; Boco et al., 2025), but even in the most metal poor systems simulated () the expected WR winds are still at least a factor of two larger than the line width observed in Hebe.
An additional problem with the WR scenario is that the fraction of WR stars decreases steeply at low metallicities (Vink and de Koter, 2005; Brinchmann et al., 2008) along with their line luminosities (Crowther and Hadfield, 2006; Crowther et al., 2023). This further suggests that WR stars are unlikely to contribute significantly at the extremely low metallicities.
Finally, the double peaked profile of the HeII emission in Hebe is not typical of WR stars, whose spectra are characterized by a Gaussian-like profile (e.g. Crowther et al., 2023), and it is not clear why the HeII WR feature should have a velocity offset relative to the H.
10 The direct collapse black hole or primordial black hole scenarios
Another interesting possibility is that Hebe is actually a very small accreting black hole. If so, given the absence of metal lines (see Rusta et al. 2026 and Übler et al. 2026 for metallicity constraints), this would be the most metal-poor accreting black hole ever discovered (see Maiolino et al., 2025, for another extreme example). Being in the vicinity of GN-z11, embedded in a strong Lyman-Werner radiation field, it would be the optimal environment for the formation of a Direct Collapse Black Hole (DCBH, e.g. Loeb and Rasio, 1994; Bromm and Loeb, 2003; Regan et al., 2017). There is no evidence for a broad component of H that would indicate the presence of a Broad Line Region. However, it could potentially be that the observed narrow H is itself tracing the BLR of a very small black hole, with very low velocity dispersion. Übler et al. (2026) infer an upper limit on the black hole mass of a few times . This would be borderline with expectations for DCBHs, but still consistent with the predictions (Ferrara et al., 2014).
Another formation channel can be that of a primordial black hole (PBH, Hawking, 1971; Carr and Hawking, 1974; Escrivà et al., 2024). These are expected to have a broad mass function, but recent constraints from the CMB suggest an upper limit of about (Matteri et al., 2025), which would more comfortably comply with the observational upper limit provided by Übler et al. (2026); although one should take into account the strong clustering expected for PBHs, which can result in a rapid growth via merging in the early Universe. In the PBH scenario, the vicinity of GN-z11 would allow the initially “naked” black holes to start accreting gas from the circumgalactic medium of the galaxy, as it plunges into its halo.
Other possible formation channels for massive black holes are stellar seeds accreting through super-Eddington bursts (e.g. Schneider et al., 2023; Trinca et al., 2023, 2024; Banik et al., 2019), or via rapid merging of stars and black holes in nuclear cluster (e.g. Partmann et al., 2025; Rantala et al., 2025). However, in these cases, the star formation associated with the black hole seeding and subsequent accretion is expected to result in significant chemical enrichment.
Regardless of the potential nature and origin, an accreting black hole in such a metal-poor environment would be an extremely interesting finding. However, there are a few aspects that make these scenarios less plausible. First, the lower limits on the EW(HeII) are uncomfortably high for the AGN scenario, as illustrated in Fig.5; even stronger tension is found by Übler et al. (2026) when exploring the EW(H). Secondly, the observed HeII line profile is quite different from the H one. While BLR stratification can broaden HeII with respect to Balmer lines (e.g. Brazzini et al., 2026), HeII in AGNs is generally not strongly blueshifted, nor it presents a double-peaked profile, unless the Balmer lines also do.
11 Excluding photoionization from the AGN in GN-z11
As mentioned in the introduction, Maiolino et al. (2024a) identified a number of spectral features indicating the presence of a type 1 AGN in GN-z11. In particular, they found evidence of gas with densities , typical of an AGN BLR, along with other AGN spectroscopic tracers, which were also identified in subsequent works (e.g., Ji et al., 2025; Scholtz et al., 2025b; Crespo Gómez et al., 2026; Fabian et al., 2026).
Hebe could then be a pristine clump of gas photoionized by the AGN in GN-z11. This scenario can be tested by adopting (and updating) the same methodology used by Maiolino et al. (2024b), i.e. by estimating the HeII luminosity expected from a cloud illuminated by an AGN-like spectrum with the luminosity estimated in GN-z11 ( erg/s), varying the distance of the cloud, and comparing it with the measured HeII line luminosity. This can be done under the conservative assumptions that the cloud has a size comparable with the size of the PSF (this is an upper limit, given that Hebe is unresolved), that it absorbs the entirety of the impinging ionizing photons from GN-z11 (while in reality the cloud could be clumpy and porous, or matter bounded, hence letting ionizing photons pass through), and that the projected distance is the actual cloud distance (in reality, it represents a lower limit). Using the several AGN models in NM22, the blue shaded region in Fig.7 shows the expected HeII luminosity of the ionized cloud as a function of distance from GN-z11 (the thickness of the region reflects the range of photoionization models), while the golden symbol shows the luminosity observed in Hebe. Photoionization from the AGN in GN-z11 fails to account for the observed HeII emission by more than two orders of magnitude (given that these are conservative upper limits).
It may still be possible that the AGN in GN-z11 was far more luminous about years (rest frame) before the epoch of the observation (corresponding to the light-travel time to the cloud), and that the HeII emission is the fossil signature of such a previous activity. However, the previous AGN luminosity of GN-z11 should have been higher than erg/s, i.e. should have been one of the most luminous quasars in the whole universe (e.g., Rakshit et al., 2020); yet, in contrast to the hyperluminous quasars, which are found at z when searching across the whole sky, GN-z11 should have been so hyperluminous already at z10, and located in the narrow Chandra Deep Field North.
Yet another possibility is that the emission of GN-z11 is highly anisotropic. In this scenario, Hebe could be seeing a radiation much larger than our line of sight. Specifically, this could be happening if GN-z11 is accreting at highly super-Eddington rate, in which case the ionizing radiation would be highly beamed in the direction perpendicular to the accretion disc (e.g. Madau and Haardt, 2024), and Hebe sees that disc nearly face on, while we are observing it close to edge on. We have explored this scenario more quantitatively by leveraging the framework presented in Madau and Maiolino (2026) for Little Red Dots and Little Blue Dots. We obtain that even in this configuration, the model falls short in explaining the HeII luminosity of Hebe by at least an order of magnitude (especially given the very conservative assumptions on the distance and property of the putative clump). Additionally, the highly anisotropic and fortuitous orientation scenario, in such a small field, would imply that there should be many other similar AGN in the sky at , pointing towards us with magnitude .
12 Summary and Conclusions
We have presented deep NIRSpec-IFU high-resolution (G235H) observations of Hebe, a HeII emitting object in the halo of the z=10.6 galaxy GN-z11, previously identified via medium resolution (G235M) IFU spectroscopy. The main results of the observations can be summarized as follows:
HeII emission is confirmed by the G235H data, in a compact region, at the location of the previous G235M detection, 3 kpc from GN-z11.
The HeII identification is independently confirmed by the detection of H in the same region by Übler et al. (2026).
No metal lines are detected.
The HeII emission is spectrally resolved in two components separated by about 120 km/s; the H emission is primarily associated with one of them.
The EW(HeII) is extremely high, i.e. Å.
From these findings we infer the following:
The only models consistent with the observed EW(HeII) and HeII/H ratios are those involving PopIII stars. Further constraints supporting this scenario come from modelling the non-detection of metal lines, as discussed in a companion paper (Rusta et al., 2026).
We have explored the possibility that the HeII emission arises from a population of extremely metal poor WR stars. This scenario is very unlikely, based on the observed properties of GN-z11 and also based on the steeply decreasing abundance and luminosity of the WR stars at low metallicities.
A Direct Collapse Black Hole or a Primordial Black Hole can potentially be invoked to explain the HeII emission. However, these scenarios face some issues that make them less plausible, although they cannot be entirely ruled out.
Photoionization by the AGN hosted in GN-z11 is safely ruled out.
In summary, Hebe represents one of the most convincing pieces of evidence for Population III stars in the early Universe, and its properties support models of their formation and early evolution. However, equally compelling alternatives, such as a Direct Collapse Black Hole or a Primordial Black Hole, cannot be completely excluded.
Acknowledgments
We are grateful to Richard Ellis, Harley Katz and Jorryt Matthee for useful comments. This work is based on observations made with the National Aeronautics and Space Administration (NASA)/European Space Agency (ESA)/Canadian Space Agency (CSA) JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the STScI, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programme PID 5086. RM, FD, JS, IJ, GJ acknowledge support from the Science and Technology Facilities Council (STFC), by the European Research Council (ERC) through Advanced Grant 695671 “QUENCH”, by the UK Research and Innovation (UKRI) Frontier Research grant RISEandFALL. RM also acknowledges support from a Royal Society Research Professorship grant. AJB, GC and JC acknowledge funding from the “FirstGalaxies” Advanced Grant from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 789056). HÜ thanks the Max Planck Society for support through the Lise Meitner Excellence Program. HÜ acknowledges funding by the European Union (ERC APEX, 101164796). MP acknowledges support through the grants PID2021-127718NB-I00, PID2024-159902NA-I00, and RYC2023-044853-I, funded by the Spain Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 and El Fondo Social Europeo Plus FSE+. JW gratefully acknowledges support from the Cosmic Dawn Center through the DAWN Fellowship. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant No. 140. PGP-G acknowledges support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia, Innovación y Universidades MCIU/AEI/10.13039/501100011033,FEDER Una manera de hacer Europa. JAAT acknowledges support from the Simons Foundation and JWST program 3215. Support for program 3215 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. BRP acknowledges support from grants PID2021-127718NB-I00 and PID2024-158856NA-I00 funded by Spanish Ministerio de Ciencia e Innovación y Universidades MCIU/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. DJE acknowledges support by JWST/NIRCam contract to the University of Arizona, NAS5-02105, and as a Simons Foundation Investigator. EE, KH and CNAW acknowledge support by the JWST/NIRCam contract to the University of Arizona NAS5-0215. SC, GV, and SZ acknowledge support by European Union’s HE ERC Starting Grant No. 101040227 - WINGS. WMB gratefully acknowledges support from DARK via the DARK fellowship. This work was supported by a research grant (VIL54489) from VILLUM FONDEN. JMH acknowledges support from JWST Program #8544. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Data availability
The NIRSpec data used in this study are publicly available at the STScI MAST archive: https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html, under GO programme 5086.
Appendix A Single Gaussian simultaneous HeII and H fit
Fig.8 shows the result of simultaneously fitting HeII and H with a single Gaussian, i.e. the the two lines are fitted with Gaussians with the same velocity and width. In this case , in contrast with the case of fitting with two Gaussian components (), resulting in a – a clear indication that the fit with a single Gaussian fit is much worse than the fit with two Gaussian components.
Appendix B Metal lines non detection
Fig.9 shows the spectrum of Hebe zoomed at the wavelengths where typically the strongest metal lines are observed in normal galaxies, specifically, from top to bottom: CIII]1906,1908, OIII]1661,1666, NIII]1749 multiplet, CIV1548,1551, and NIV]1483,1486. Note that an unmasked faint open shutter/artefact affects a region near the NIV doublet (gray shaded region).
Appendix C Gravitational lens modeling
Both GNz-11 and Hebe are in close proximity (i.e., within projected arcsecond) of a foreground galaxy at . In order to determine the lensing magnification of each object, we use the code pyautolens (Nightingale et al., 2021) to construct the gravitational lens model of this field.
First, we collapse our data cube around the redshifted wavelength of H of the foreground galaxy (), resulting in a map containing a strong, resolved detection of the foreground galaxy as well as a continuum detection of GNz-11. We then create a model of this field by using Sérsic light profiles for the two galaxies, and use pyautolens to determine the best-fit properties of each. The mass of the foreground galaxy is assumed to be an isothermal profile with the same spatial centroid, axis ratio, and position angle as the best-fit Sérsic profile. The Einstein radius of this mass profile is determined assuming a total (i.e., baryonic and non-baryonic) mass of , as derived through kinematic modeling of this source.
This approach results in gravitational magnification estimates of for Hebe and for GNz-11. These values are dependent on the total mass of the foreground galaxy, so there is some uncertainty in each. If we instead adopt the conservative total mass limit of Tacchella et al. (2023, ), which was determined for the total halo mass containing GNz-11, then we find slightly higher magnification for both Hebe () and GNz-11 (). If we instead adopt the best-fit stellar mass based on SED fits (), then we find that each galaxy is only magnified by a few percent. In any case, we may rule out very high magnification factors of for each source.